If you have a 86 what will you need to have a 90 average?
Found 2 solutions by strawberrygirl183, rfer:
Answer by strawberrygirl183(35)

4 more points

Answer by rfer(16322)

(x+86)/2=90
x+86=90*2
x+86=180
x+180-86
x=94